OpenStudy (anonymous):

integral of (x)Ln(x)/(sqrtx^2-1)

OpenStudy (anonymous):

i think im gonna try integration by parts

OpenStudy (anonymous):

anyone there?

zepdrix (zepdrix):

\[\large \int\limits \frac{x \ln x}{\sqrt{x^2-1}}dx \qquad=\qquad \frac{1}{2}\int\limits \ln x \left(\frac{2x}{\sqrt{x^2-1}}dx\right)\] Yah by parts sounds correct. I grouped this a little strange so you can easily see your u and dv. Lemme know if that's still a bit confusing.

zepdrix (zepdrix):

Oh woops I didn't need to add the 1/2 and 2 .. my mistake.. got my derivatives mixed up. \[\large \int\limits\limits \frac{x \ln x}{\sqrt{x^2-1}}dx \qquad=\qquad \int\limits\limits \ln x \left(\frac{x}{\sqrt{x^2-1}}dx\right)\]

OpenStudy (anonymous):

i figured it out. I had to use integration by parts